The Ancient Greeks thought that the most pleasing shape for a rectangle was one for which the ratio of the length to the width was 8 to 5, the golden ratio. If the length of a rectangle painting is 2 feet longer than its width, then for what dimensions would the length and width have the golden ratio?
Answer by checkley71(8403)

8/5=(X+2)/X CROSS MULTIPLY
8X=5(X+2)
8X=5X+10
8X-5X=10
3X=10
X=10/3 ANSWER FOR THE WIDTH.
10/3+2=(10+6)/3=16/3 FOR THE LENGTH
PROOF
8/5=(16/3)*(3/10) HERE WE INVERT THE DENOMINATOR (10/3) & MULTIPLY.
8/5=16/10
8/5=8/5